OpenStudy (anonymous):

whats the area of a rectangle if the side is 1 and 2/3 and the length is 2/3

OpenStudy (anonymous):

multiply \[1\tfrac{2}{3}\times \frac{2}{3}\]

OpenStudy (anonymous):

\[1\frac{ 2 }{ 3 }=\frac{ 3+2 }{ 3 }=\frac{ 5 }{ 3 }\] Area of rectangle = l*b = 5/3 *2/3=10/9 sq units.

OpenStudy (anonymous):

need to start with \[\frac{5}{3}\times \frac{2}{3}\] then multiply straight across

OpenStudy (anonymous):

yes i did that and it came up as 10/9 but i need to put that into an actual number because the answers are not fractions

OpenStudy (anonymous):

10/9 = 1.11 sq units